14. United Sales Association (USA) recently borrowed $25,000 from its bank at a simple interest rate of 11 percent. The loan is for one year, and the loan agreement calls for the interest to be added to the amount borrowed and the total amount to be repaid in monthly installments. (a) What are the loan’s monthly payments? (b) Compute the loan’s APR and, rEAR.
Answer #1

ANSWER:

(a)

Simple interest=(Principal*Interest rate*Time period)

=($25000*11%*1)=$2750

Hence total amount payable=SI+Principal

=(25000+2750)=$27,750

Hence loan's monthly payments are =($27,750/12)

=$2312.5

---------------------------------------------------------------

As HOMEWORKLIB RULES rule we answered (a) only. repost (b) question

_____________________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU**************